# ADOPT: Modified Adam Can Converge with Any $\beta_{2}$ with the Optimal Rate

### v1

@@ Appendix F Proofs
$\leq\frac{f_{\mathrm{sup}}-f_{\mathrm{inf}}}{\alpha}T+\frac{\alpha\left(1+\left(2\sqrt{2}-1\right)\beta_{1}\right)G^{2}LT}{2\left(1-\beta_{1}\right)\epsilon^{2}}+\frac{\sqrt{2}\beta_{1}G^{2}}{\left(1-\beta_{1}\right)\epsilon}\left(2\sqrt{T}-1\right)$
$\quad\ +\frac{\alpha\beta_{1}\left(1+\left(2\sqrt{2}-1\right)\beta_{1}\right)G^{2}L}{\sqrt{2}\left(1-\beta_{1}\right)^{2}\epsilon^{2}}\left(1+\log T\right)+\frac{\alpha\beta_{1}^{2}G^{2}L}{\left(1-\beta_{1}\right)^{2}\epsilon^{2}}\left(2-\frac{1}{T}\right)$ (47)
@@
$\leq\frac{3C_{T}\left(f_{\mathrm{sup}}-f_{\mathrm{inf}}\right)}{2\alpha}\frac{1}{\sqrt{T}}+\frac{3\alpha\left(1+\left(2\sqrt{2}-1\right)\beta_{1}\right)C_{T}G^{2}L}{4\left(1-\beta_{1}\right)\epsilon^{2}\sqrt{T}}+\frac{3\beta_{1}C_{T}G^{2}}{\sqrt{2}\left(1-\beta_{1}\right)\epsilon}\left(\frac{2}{T}-\frac{1}{T^{3/2}}\right)$
$\quad\ +\frac{3\alpha\beta_{1}\left(1+\left(2\sqrt{2}-1\right)\beta_{1}\right)C_{T}G^{2}L}{2\sqrt{2}\left(1-\beta\right)^{2}\epsilon^{2}}\left(\frac{1}{T^{3/2}}+\frac{\log T}{T^{3/2}}\right)+\frac{3\alpha\beta_{1}^{2}C_{T}G^{2}L}{2\left(1-\beta_{1}\right)^{2}\epsilon^{2}}\left(\frac{2}{T^{3/2}}-\frac{1}{T^{5/2}}\right),$ (50)
where $C_{T}=\sqrt{\left(1-\beta_{2}^{T}\right)G^{2}+\epsilon^{2}}$.
@@ Appendix G Lemmas
For the ADOPT algorithm, the following holds for $t\geq 1$:
${\bm{\phi}}_{t}-{\bm{\phi}}_{t-1}$ $=\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}}{1-\beta_{1}}{\bm{m}}_{t-1}-\alpha_{t}\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},$ (95)
@@
$=-\frac{\alpha_{1}{\bm{g}}_{1}}{\max\left\{\sqrt{{\bm{v}}_{0}},\epsilon\right\}}.$ (98)
@@
${\bm{\phi}}_{t}-{\bm{\phi}}_{t-1}$ $=\frac{1}{1-\beta_{1}}\left({\bm{\theta}}_{t}-{\bm{\theta}}_{t-1}\right)-\frac{\beta_{1}}{1-\beta_{1}}\left({\bm{\theta}}_{t-1}-{\bm{\theta}}_{t-2}\right)$ (99)
@@
$=\frac{1}{1-\beta_{1}}\left(\alpha_{t-1}\beta_{1}{\bm{m}}_{t-1}-\alpha_{t}\left(\beta_{1}{\bm{m}}_{t-1}+\left(1-\beta_{1}\right)\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}}\right)\right)$ (101)
$=\frac{1}{1-\beta_{1}}\left(\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}{\bm{m}}_{t-1}-\alpha_{t}\left(1-\beta_{1}\right)\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}}\right)$ (102)
$=\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}}{1-\beta_{1}}{\bm{m}}_{t-1}-\alpha_{t}\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}}$ (103)
@@
For the ADOPT algorithm, the following holds for $t\geq 1$:
@@
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}\left(1-\beta_{1}^{t-1}\right)G^{2}}{\left(1-\beta_{1}\right)\sqrt{\beta_{2}^{t-2}+\epsilon^{2}}}-\alpha_{t}\frac{\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|_{i}^{4/3}\right]^{3/2}}{\sqrt{\left(1-\beta_{2}^{t}\right)G^{2}+\epsilon^{2}}}.$ (108)
@@
$=\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}}{1-\beta_{1}}\nabla f\left({\bm{\theta}}_{t-1}\right))^{\top}{\bm{m}}_{t-1}-\alpha_{t}\nabla f\left({\bm{\theta}}_{t-1}\right)^{\top}\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}}$ (109)
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}}{1-\beta_{1}}\left\|\nabla f\left({\bm{\theta}}_{t-1}\right))\right\|\left\|{\bm{m}}_{t-1}\right\|-\alpha_{t}\nabla f\left({\bm{\theta}}_{t-1}\right)^{\top}\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}}$ (110)
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}G}{1-\beta_{1}}\left\|{\bm{m}}_{t-1}\right\|-\alpha_{t}\nabla f\left({\bm{\theta}}_{t-1}\right)^{\top}\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}}.$ (111)
@@
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}G}{1-\beta_{1}}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|\right]-\alpha_{t}\mathbb{E}\left[\nabla f\left({\bm{\theta}}_{t-1}\right)^{\top}\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}}\right]$ (112)
@@
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}G}{1-\beta_{1}}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|\right]-\alpha_{t}\sum_{i=1}^{D}\mathbb{E}\left[\frac{\left(\nabla f\left({\bm{\theta}}_{t-1}\right)\right)_{i}\cdot\left({\bm{g}}_{t}\right)_{i}}{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}\right]$ (113)
@@
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}G}{1-\beta_{1}}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|\right]-\alpha_{t}\sum_{i=1}^{D}\mathbb{E}\left[\frac{\left(\nabla f\left({\bm{\theta}}_{t-1}\right)\right)_{i}^{2}}{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}\right]$ (114)
@@
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}G}{1-\beta_{1}}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|\right]-\alpha_{t}\sum_{i=1}^{D}\mathbb{E}\left[\frac{\left(\nabla f\left({\bm{\theta}}_{t-1}\right)\right)_{i}^{2}}{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}+\epsilon^{2}}}\right]$ (115)
@@
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}G}{1-\beta_{1}}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|\right]-\alpha_{t}\mathbb{E}\left[\frac{\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|^{2}}{\sqrt{\sum_{i=1}^{D}\left({\bm{v}}_{t-1}\right)_{i}+\epsilon^{2}}}\right]$ (116)
@@
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}G}{1-\beta_{1}}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|\right]-\alpha_{t}\frac{\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|_{i}^{4/3}\right]^{3/2}}{\sqrt{\mathbb{E}\left[\sum_{i=1}^{D}\left({\bm{v}}_{t-1}\right)_{i}\right]+\epsilon^{2}}}$ (117)
@@
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}G}{1-\beta_{1}}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|\right]-\alpha_{t}\frac{\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|_{i}^{4/3}\right]^{3/2}}{\sqrt{\left(1-\beta_{2}^{t}\right)G^{2}+\epsilon^{2}}}$ (118)
@@
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}\left(1-\beta_{1}^{t-1}\right)G^{2}}{\left(1-\beta_{1}\right)\sqrt{\beta_{2}^{t-2}+\epsilon^{2}}}-\alpha_{t}\frac{\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|_{i}^{4/3}\right]^{3/2}}{\sqrt{\left(1-\beta_{2}^{t}\right)G^{2}+\epsilon^{2}}}.$ (119)
@@
$\mathbb{E}\left[\sum_{i=1}^{D}\left({\bm{v}}_{t}\right)_{i}\right]\leq\left(1-\beta_{2}^{t}\right)G^{2}.$ (120)

### v3

@@ Appendix F Proofs
$\leq\frac{f_{\mathrm{sup}}-f_{\mathrm{inf}}}{\alpha}T+\frac{\alpha\left(1+\left(2\sqrt{2}-1\right)\beta_{1}\right)G^{2}LT}{\left(1-\beta_{1}\right)\epsilon^{2}}+\frac{G^{4}T}{2c^{2}\epsilon^{3}}+\frac{{2}\beta_{1}G^{2}}{\left(1-\beta_{1}\right)\epsilon}\left(2\sqrt{T}-1\right)$
$\quad\ +\frac{\sqrt{2}\alpha\beta_{1}\left(1+\left(2\sqrt{2}-1\right)\beta_{1}\right)G^{2}L}{\left(1-\beta_{1}\right)^{2}\epsilon^{2}}\left(1+\log T\right)+\frac{2\alpha\beta_{1}^{2}G^{2}L}{\left(1-\beta_{1}\right)^{2}\epsilon^{2}}\left(2-\frac{1}{T}\right)$ (51)
@@
$\leq\frac{3C\left(f_{\mathrm{sup}}-f_{\mathrm{inf}}\right)}{2\alpha}\frac{1}{\sqrt{T}}+\frac{3\alpha\left(1+\left(2\sqrt{2}-1\right)\beta_{1}\right)CG^{2}L}{2\left(1-\beta_{1}\right)\epsilon^{2}\sqrt{T}}+\frac{3G^{4}T}{4c^{2}\epsilon^{3}}+\frac{3\beta_{1}CG^{2}}{\left(1-\beta_{1}\right)\epsilon}\left(\frac{2}{T}-\frac{1}{T^{3/2}}\right)$
$\quad\ +\frac{3\alpha\beta_{1}\left(1+\left(2\sqrt{2}-1\right)\beta_{1}\right)CG^{2}L}{\sqrt{2}\left(1-\beta\right)^{2}\epsilon^{2}}\left(\frac{1}{T^{3/2}}+\frac{\log T}{T^{3/2}}\right)+\frac{3\alpha\beta_{1}^{2}CG^{2}L}{\left(1-\beta_{1}\right)^{2}\epsilon^{2}}\left(\frac{2}{T^{3/2}}-\frac{1}{T^{5/2}}\right)$ (54)
$=\mathcal{O}\left(1/\sqrt{T}\right),$ (55)
where $C=2\sqrt{G^{2}+\epsilon^{2}}$.
@@ Appendix G Lemmas
For the clipped ADOPT algorithm, the following holds for $t\geq 1$:
${\bm{\phi}}_{t}-{\bm{\phi}}_{t-1}$ $=\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}}{1-\beta_{1}}{\bm{m}}_{t-1}-\alpha_{t}\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right),$ (100)
@@
$=-\alpha_{1}\cdot\mathrm{Clip}\left(\frac{{\bm{g}}_{1}}{\max\left\{\sqrt{{\bm{v}}_{0}},\epsilon\right\}},c_{t}\right).$ (103)
@@
${\bm{\phi}}_{t}-{\bm{\phi}}_{t-1}$ (104)
$=\frac{1}{1-\beta_{1}}\left({\bm{\theta}}_{t}-{\bm{\theta}}_{t-1}\right)-\frac{\beta_{1}}{1-\beta_{1}}\left({\bm{\theta}}_{t-1}-{\bm{\theta}}_{t-2}\right)$ (105)
@@
$=\frac{1}{1-\beta_{1}}\left(\alpha_{t-1}\beta_{1}{\bm{m}}_{t-1}-\alpha_{t}\left(\beta_{1}{\bm{m}}_{t-1}+\left(1-\beta_{1}\right)\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)\right)\right)$ (107)
$=\frac{1}{1-\beta_{1}}\left(\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}{\bm{m}}_{t-1}-\alpha_{t}\left(1-\beta_{1}\right)\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)\right)$ (108)
$=\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}}{1-\beta_{1}}{\bm{m}}_{t-1}-\alpha_{t}\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)$ (109)
@@
For the clipped ADOPT algorithm, the following holds for $t\geq 1$:
@@
$\leq\frac{\sqrt{2}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}\left(1-\beta_{1}^{t-1}\right)G^{2}}{\left(1-\beta_{1}\right)\epsilon}-\frac{\alpha_{t}}{2}\frac{\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|_{i}^{4/3}\right]^{3/2}}{\sqrt{G^{2}+\epsilon^{2}}}+\frac{\alpha_{t}G^{4}}{2\epsilon^{3}c_{t}^{2}}.$ (114)
@@
$=\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}}{1-\beta_{1}}\nabla f\left({\bm{\theta}}_{t-1}\right))^{\top}{\bm{m}}_{t-1}-\alpha_{t}\nabla f\left({\bm{\theta}}_{t-1}\right)^{\top}\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)$ (115)
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}}{1-\beta_{1}}\left\|\nabla f\left({\bm{\theta}}_{t-1}\right))\right\|\left\|{\bm{m}}_{t-1}\right\|-\alpha_{t}\nabla f\left({\bm{\theta}}_{t-1}\right)^{\top}\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)$ (116)
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}G}{1-\beta_{1}}\left\|{\bm{m}}_{t-1}\right\|-\alpha_{t}\nabla f\left({\bm{\theta}}_{t-1}\right)^{\top}\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right).$ (117)
@@
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}G}{1-\beta_{1}}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|\right]-\alpha_{t}\mathbb{E}\left[\nabla f\left({\bm{\theta}}_{t-1}\right)^{\top}\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)\right]$ (118)
@@
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}G}{1-\beta_{1}}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|\right]-\frac{\alpha_{t}}{2}\sum_{i=1}^{D}\mathbb{E}\left[\frac{\left(\nabla f\left({\bm{\theta}}_{t-1}\right)\right)_{i}\cdot\left({\bm{g}}_{t}\right)_{i}}{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}\right]+\frac{\alpha_{t}G^{4}}{2\epsilon^{3}c_{t}^{2}}$ (119)
@@
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}G}{1-\beta_{1}}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|\right]-\frac{\alpha_{t}}{2}\sum_{i=1}^{D}\mathbb{E}\left[\frac{\left(\nabla f\left({\bm{\theta}}_{t-1}\right)\right)_{i}^{2}}{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}\right]+\frac{\alpha_{t}G^{4}}{2\epsilon^{3}c_{t}^{2}}$ (120)
@@
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}G}{1-\beta_{1}}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|\right]-\frac{\alpha_{t}}{2}\sum_{i=1}^{D}\mathbb{E}\left[\frac{\left(\nabla f\left({\bm{\theta}}_{t-1}\right)\right)_{i}^{2}}{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}+\epsilon^{2}}}\right]+\frac{\alpha_{t}G^{4}}{2\epsilon^{3}c_{t}^{2}}$ (121)
@@
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}G}{1-\beta_{1}}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|\right]-\frac{\alpha_{t}}{2}\mathbb{E}\left[\frac{\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|^{2}}{\sqrt{\sum_{i=1}^{D}\left({\bm{v}}_{t-1}\right)_{i}+\epsilon^{2}}}\right]+\frac{\alpha_{t}G^{4}}{2\epsilon^{3}c_{t}^{2}}$ (122)
@@
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}G}{1-\beta_{1}}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|\right]-\frac{\alpha_{t}}{2}\frac{\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|_{i}^{4/3}\right]^{3/2}}{\sqrt{\mathbb{E}\left[\sum_{i=1}^{D}\left({\bm{v}}_{t-1}\right)_{i}\right]+\epsilon^{2}}}+\frac{\alpha_{t}G^{4}}{2\epsilon^{3}c_{t}^{2}}$ (123)
@@
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}G}{1-\beta_{1}}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|\right]-\frac{\alpha_{t}}{2}\frac{\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|_{i}^{4/3}\right]^{3/2}}{\sqrt{G^{2}+\epsilon^{2}}}+\frac{\alpha_{t}G^{4}}{2\epsilon^{3}c_{t}^{2}}$ (124)
@@
$\leq\frac{\sqrt{2}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}\left(1-\beta_{1}^{t-1}\right)G^{2}}{\left(1-\beta_{1}\right)\epsilon}-\frac{\alpha_{t}}{2}\frac{\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|_{i}^{4/3}\right]^{3/2}}{\sqrt{G^{2}+\epsilon^{2}}}+\frac{\alpha_{t}G^{4}}{2\epsilon^{3}c_{t}^{2}}.$ (125)
Lemma G.10.
For the clipped ADOPT algorithm, the following holds for $t\geq 0$:
$2\nabla f\left({\bm{\theta}}_{t-1}\right)^{\top}\mathbb{E}\left[\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)\right]$ (126)
$\geq\sum_{i=1}^{D}\frac{\left(\nabla f\left({\bm{\theta}}_{t-1}\right)\right)_{i}^{2}}{{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}}-\frac{G^{4}}{\epsilon^{3}c_{t}^{2}}$ (127)
Proof.
$\mathbb{E}\left[\mathrm{Clip}\left(\frac{\left({\bm{g}}_{t}\right)_{i}}{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}},c_{t}\right)\right]$ (128)
$=\mathbb{E}\left[\left(\delta_{t,i}\cdot\frac{c_{t}}{\left|\left({\bm{g}}_{t}\right)_{i}/{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}\right|}+\left(1-\delta_{t,i}\right)\right)\cdot\frac{\left({\bm{g}}_{t}\right)_{i}}{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}\right]$ (129)
$=\frac{\left(\nabla f\left({\bm{\theta}}_{t-1}\right)\right)_{i}}{{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}}-\mathbb{E}\left[\delta_{t,i}\left(1-\frac{c_{t}}{\left|\left({\bm{g}}_{t}\right)_{i}/{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}\right|}\right)\cdot\frac{\left({\bm{g}}_{t}\right)_{i}}{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}\right],$ (130)
where $\delta_{t,i}$ is an indicator function of whether the normalized gradient is clipped:
$\delta_{t,i}={\bm{1}}_{\left\{\left|{\left({\bm{g}}_{t}\right)_{i}}/{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}\right|>c_{t}\right\}}.$ (131)
Using this equation, the following inequality is derived:
$\left(\frac{\left(\nabla f\left({\bm{\theta}}_{t-1}\right)\right)_{i}}{{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}}-\mathbb{E}\left[\mathrm{Clip}\left(\frac{\left({\bm{g}}_{t}\right)_{i}}{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}},c_{t}\right)\right]\right)^{2}$ (132)
$=\mathbb{E}\left[\delta_{t,i}\left(1-\frac{c_{t}}{\left|\left({\bm{g}}_{t}\right)_{i}/{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}\right|}\right)\cdot\frac{\left({\bm{g}}_{t}\right)_{i}}{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}\right]^{2}$ (133)
$=\mathbb{E}\left[\delta_{t,i}\right]^{2}\mathbb{E}\left[\left(1-\frac{c_{t}}{\left|\left({\bm{g}}_{t}\right)_{i}/{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}\right|}\right)\cdot\frac{\left({\bm{g}}_{t}\right)_{i}}{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}\mid\delta_{t,i}=1\right]^{2}$ (134)
$\leq\mathbb{E}\left[\delta_{t,i}\right]^{2}\mathbb{E}\left[\left(1-\frac{c_{t}}{\left|\left({\bm{g}}_{t}\right)_{i}/{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}\right|}\right)^{2}\cdot\frac{\left({\bm{g}}_{t}\right)_{i}^{2}}{\max\left\{\left({\bm{v}}_{t-1}\right)_{i},\epsilon^{2}\right\}}\mid\delta_{t,i}=1\right]$ (135)
$\leq\mathbb{E}\left[\delta_{t,i}\right]^{2}\mathbb{E}\left[\frac{\left({\bm{g}}_{t}\right)_{i}^{2}}{\max\left\{\left({\bm{v}}_{t-1}\right)_{i},\epsilon^{2}\right\}}\mid\delta_{t,i}=1\right]$ (136)
$=\mathbb{E}\left[\delta_{t,i}\right]\mathbb{E}\left[\frac{\left({\bm{g}}_{t}\right)_{i}^{2}}{\max\left\{\left({\bm{v}}_{t-1}\right)_{i},\epsilon^{2}\right\}}\right]$ (137)
$=\mathrm{Pr}\left[\left|\frac{\left({\bm{g}}_{t}\right)_{i}}{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}\right|>c_{t}\right]\mathbb{E}\left[\frac{\left({\bm{g}}_{t}\right)_{i}^{2}}{\max\left\{\left({\bm{v}}_{t-1}\right)_{i},\epsilon^{2}\right\}}\right]$ (138)
$\leq\frac{1}{c_{t}^{2}}\mathbb{E}\left[\frac{\left({\bm{g}}_{t}\right)_{i}^{2}}{\max\left\{\left({\bm{v}}_{t-1}\right)_{i},\epsilon^{2}\right\}}\right]^{2},$ (139)
where the first inequality is due to the Jensen’s inequality, and the last inequality is due to the Chebyshev’s inequality. Therefore, the following holds:
$2\left(\nabla f\left({\bm{\theta}}_{t-1}\right)\right)_{i}\cdot\mathbb{E}\left[\mathrm{Clip}\left(\frac{\left({\bm{g}}_{t}\right)_{i}}{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}},c_{t}\right)\right]$ (140)
$=\frac{\left(\nabla f\left({\bm{\theta}}_{t-1}\right)\right)_{i}^{2}}{{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}}+{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}\mathbb{E}\left[\mathrm{Clip}\left(\frac{\left({\bm{g}}_{t}\right)_{i}}{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}},c_{t}\right)\right]^{2}$ (141)
$\quad-{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}\left(\frac{\left(\nabla f\left({\bm{\theta}}_{t-1}\right)\right)_{i}}{{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}}-\mathbb{E}\left[\mathrm{Clip}\left(\frac{\left({\bm{g}}_{t}\right)_{i}}{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}},c_{t}\right)\right]\right)^{2}$
$\geq\frac{\left(\nabla f\left({\bm{\theta}}_{t-1}\right)\right)_{i}^{2}}{{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}}$ (142)
$\quad-{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}\left(\frac{\left(\nabla f\left({\bm{\theta}}_{t-1}\right)\right)_{i}}{{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}}-\mathbb{E}\left[\mathrm{Clip}\left(\frac{\left({\bm{g}}_{t}\right)_{i}}{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}},c_{t}\right)\right]\right)^{2}$
$\geq\frac{\left(\nabla f\left({\bm{\theta}}_{t-1}\right)\right)_{i}^{2}}{{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}}-\frac{1}{c_{t}^{2}\left(\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}\right)^{3}}\mathbb{E}\left[\left({\bm{g}}_{t}\right)_{i}^{2}\right]^{2}$ (143)
$\geq\frac{\left(\nabla f\left({\bm{\theta}}_{t-1}\right)\right)_{i}^{2}}{{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}}-\frac{1}{\epsilon^{3}c_{t}^{2}}\mathbb{E}\left[\left({\bm{g}}_{t}\right)_{i}^{2}\right]^{2}$ (144)
By summing it up for all the dimensions, the following inequality is derived.
$2\nabla f\left({\bm{\theta}}_{t-1}\right)^{\top}\mathbb{E}\left[\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)\right]$ (145)
$\geq\sum_{i=1}^{D}\frac{\left(\nabla f\left({\bm{\theta}}_{t-1}\right)\right)_{i}^{2}}{{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}}-\frac{1}{\epsilon^{3}c_{t}^{2}}\mathbb{E}\left[\left({\bm{g}}_{t}\right)_{i}^{2}\right]^{2}$ (146)
$\geq\sum_{i=1}^{D}\frac{\left(\nabla f\left({\bm{\theta}}_{t-1}\right)\right)_{i}^{2}}{{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}}-\frac{1}{\epsilon^{3}c_{t}^{2}}\mathbb{E}\left[\left\|{\bm{g}}_{t}\right\|^{2}\right]^{2}$ (147)
$\geq\sum_{i=1}^{D}\frac{\left(\nabla f\left({\bm{\theta}}_{t-1}\right)\right)_{i}^{2}}{{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}}-\frac{G^{4}}{\epsilon^{3}c_{t}^{2}}$ (148)
@@
$\mathbb{E}\left[\sum_{i=1}^{D}\left({\bm{v}}_{t}\right)_{i}\right]\leq G^{2}.$ (149)